OpenStudy (anonymous):

Write -0.0000356 in scientific notation ... help

OpenStudy (kirbykirby):

\(\large -3.56 \times 10^{-5}\)

OpenStudy (kirbykirby):

you need a digit between 1 and 9 before the decimal, then to find the exponent... since you have a number less than 1, then the exponent is negative, and the "value" of the exponent is the number of zeros after the decimal + 1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

Write this number without exponents \[5.867\times10^{-4}\]

OpenStudy (kirbykirby):

Think of the reverse process... you have a negative exponent so you have a number less than 1, so something like 0. ____ Now, the exponent is 4, so that means you have 3 zeros after the decimal... so 0.0005867

OpenStudy (anonymous):

thats how you convert it ?

OpenStudy (anonymous):

thanks

OpenStudy (kirbykirby):

yeah

OpenStudy (kirbykirby):

:)

OpenStudy (anonymous):

write the answer without exponents \[(5x10^{3}) \times (6x10^{4})\]

OpenStudy (kirbykirby):

first, manipulate it with the exponents: \((5 \times 10^3) \times (6 \times 10^4)=5\times 6 \times 10^3 \times 10^4=30\times \color{red}{10^7}=30\color{red}{0,000,000}\), here you just add the number of zeros corresponding to the exponent you have in your base 10

OpenStudy (kirbykirby):

(since the exponent is positive)

OpenStudy (anonymous):

what would be the answer then ? i dont get it

OpenStudy (anonymous):

300,000,000

OpenStudy (kirbykirby):

yes, I wrote it up there too

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Divide and write the result using scientific notation. \[\frac{ 4\times10^{8} }{ 2\times10^{-5} }\]

OpenStudy (anonymous):

how do you do this one ?

OpenStudy (kirbykirby):

divide the 4 and the 2 together, and the base 10's together using the exponent laws

OpenStudy (anonymous):

itll be 2 .. which exponent law do i use ?

OpenStudy (kirbykirby):

\[ \large \frac{a^x}{a^y}=a^{x-y}\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

whats the next step ?

OpenStudy (anonymous):

?

OpenStudy (kirbykirby):

you multiply it out.. here i'm saying \(a = 10\) and \(x = 8\) and \(y = -5\)

OpenStudy (kirbykirby):

well you don't have to fully multiply it out because it will be already in scientific notation :)

OpenStudy (anonymous):

\[\frac{ 10^{8} }{ 10^{-5} }=10^{8-5}\]

OpenStudy (kirbykirby):

\(10^{8-(-5)}\)

OpenStudy (kirbykirby):

= \(10^{13}\)

OpenStudy (anonymous):

i have to solve for the parenth rite ?

OpenStudy (anonymous):

or just leave it like that ?

OpenStudy (kirbykirby):

well it just become \(2 \times 10^{13}\) :)